OpenStudy (megannicole51):

Split into partial fractions.... 5(1+s)/s(s^2+6s+5)

OpenStudy (megannicole51):

so i know the denominator factors into s(s+3)(s+2)

OpenStudy (anonymous):

good what is next ?

OpenStudy (megannicole51):

i have to split it up in 3 different fractions?

OpenStudy (megannicole51):

A, B, and C?

OpenStudy (anonymous):

yes

OpenStudy (megannicole51):

so ill have A/s + B/(s+3) + C/(s+2) ?

OpenStudy (anonymous):

yes!

OpenStudy (megannicole51):

yay!

OpenStudy (megannicole51):

so this is where i get mixed up so hold on lol

OpenStudy (megannicole51):

so then i set A+B+C=5(1+s)?

OpenStudy (megannicole51):

or cross multiply first?

OpenStudy (anonymous):

now you should multiply both sides by s(s+3)(s+2)

OpenStudy (dan815):

we gotta get common denominator right

OpenStudy (dan815):

how can A+b +c equal that otherwise

OpenStudy (megannicole51):

so like A (s(s+3)(s+2))?

OpenStudy (anonymous):

A(s(s+3)(s+2))/s + B(s(s+3)(s+2))/(s+3) + C(s(s+3)(s+2))/(s+2) = 5(1+s)

OpenStudy (dan815):